OpenStudy (anonymous):

What is the value of the expression? 5^-6*5^2 ---------- 5^-8 A. 5^−20 B. 5^−12 C. 5^4 D. 5^-4

Nnesha (nnesha):

exponent rule \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\] there is negative exponent in the denominator so move that to the top

OpenStudy (anonymous):

....

Nnesha (nnesha):

sorry i don't understand dot language.

OpenStudy (anonymous):

lol sorry It's just I don't get it :'(

Nnesha (nnesha):

basses are same so just transfer -8 to the top here is an example \[\large\rm \frac{ 2^3 }{ 2^{-5} } = x^{2\color{red}{+}5}\]

OpenStudy (anonymous):

ok

Nnesha (nnesha):

and then apply exponent rule \[\huge\rm x^m \times x^n = x^{m+n}\]when you multiply same bases you should add their exponents

OpenStudy (anonymous):

ok

Nnesha (nnesha):

so what did you get ?

OpenStudy (anonymous):

7?

Nnesha (nnesha):

how did you get that ?

OpenStudy (anonymous):

2+5

OpenStudy (anonymous):

im so wrong! I'm sorry i'm being stupid

Nnesha (nnesha):

i know ... reread what i said above then try to solve

OpenStudy (anonymous):

ill just put d bye

Nnesha (nnesha):

you should familiar with the exponent rules you can't have the negative exponent there \[\huge\rm x^{-m}=\frac{ 1 }{ x^m }\]

OpenStudy (anonymous):

???

Nnesha (nnesha):

\[\huge\rm \frac{ \color{Red}{5^{-6} \times 5^{2} }}{ 5^{-8}}= \] first look at the numerator when you multiply same bases you should ADD their exponents \[\large\rm x^m \times x^n = x^{m+n}\]

OpenStudy (anonymous):

8?

Nnesha (nnesha):

-6+(2) = 8 ?

OpenStudy (anonymous):

-4

Nnesha (nnesha):

yes right \[\huge\rm \frac{ 5^{-4} }{ 5^{-8}}\] now move the -8 exponent to the top here is an example \[\huge\rm \frac{ x^m }{ x^{n} } = x^{m+(-n)}\]

OpenStudy (anonymous):

i divide now?

Nnesha (nnesha):

now you don't divide bases are the same right 5 is base and there is negative -8 exponent in the denominator move that to top of the fraction

OpenStudy (anonymous):

huh

Nnesha (nnesha):

here is an example \[\huge\rm \frac{ 2^{-3} }{ 2^{-9} } = 2^{-3+(-9)}\]

OpenStudy (anonymous):

Add -3 and -9? -12

Nnesha (nnesha):

that's an example and there is typo wait a sec

OpenStudy (anonymous):

oh

Nnesha (nnesha):

here is an example \[\huge\rm \frac{ 2^{-3} }{ 2^{-9} } = 2^{-3-(-9)}\]

OpenStudy (anonymous):

oh subtract?

OpenStudy (anonymous):

its 6

Nnesha (nnesha):

yes right - times -9 is = positive 9 so -3 +9 now look at ur question \[\huge\rm \frac{ 5^{-4} }{ 5^{-8} }= 5^{???}\]

OpenStudy (anonymous):

-12

Nnesha (nnesha):

no remember sign would change

Nnesha (nnesha):

\[\huge\rm \frac{ 5^{-4} }{ 5^{-8} }= 5^{-4-(-8)}\]

OpenStudy (anonymous):

6?

Nnesha (nnesha):

how did you get 6 ?? :O

OpenStudy (anonymous):

IDK

Nnesha (nnesha):

so why did you say 6 ?? just add the exponent that's it

OpenStudy (anonymous):

4

Nnesha (nnesha):

\[\huge\rm \frac{ 5^{-4} }{ 5^{-8} }= 5^{\color{Red}{-4-(-8)}}\] know how do solve red part(exponents ?

Nnesha (nnesha):

yes right!!

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

???

Nnesha (nnesha):

ye...

OpenStudy (anonymous):

tysm!!!

Nnesha (nnesha):

YW